# Shanshan Zhong1, Hefeng Wu1, Liang Lin1

### v1

@@ Abstract
Routing large language models (LLMs) is a novel paradigm that recommends the most suitable LLM from a pool of candidates to process a given input through a well-designed router. Our comprehensive analysis reveals a model-level scaling up phenomenon in LLMs, i.e., a capable router can significantly enhance the performance of this paradigm as the number of candidates increases. This improvement can even easily surpass the performance of the best single model in the pool and most existing strong LLMs, making it a highly promising paradigm. However, the lack of comprehensive and open-source benchmarks for Routing LLMs has hindered the development of routers. In this paper, we introduce RouterEval, a benchmark designed specifically for router research, which includes over 200,000,000 performance records for 12 popular LLM evaluations across areas such as knowledge-based Q&A, commonsense reasoning, semantic understanding, mathematical reasoning, and instruction following, based on more than 8,500 LLMs. Using RouterEval, extensive evaluations of existing Routing LLM methods reveal that most still have significant room for improvement. See project page for all data, code and tutorial .
@@
Zhongzhan Huang1, Guoming Ling1, Vincent S. Liang2, Yupei Lin1, Yandong Chen1, Shanshan Zhong1, Hefeng Wu1, Liang Lin1 1Sun Yat-sen University, 2Purdue University
@@ 1 Introduction
Routing LLMs Jitkrittum et al. (2025); Lu et al. (2023a); Zhao et al. (2024a); Shnitzer et al. (2023a) aims to establish an efficient router capable of selecting an LLM from a pool of candidates with varying abilities to effectively handle a given input, in order to achieve specific goals such as high overall accuracy, low computational cost, or minimal hallucination, as shown in Fig. 1. Since this paradigm only involves input assignment, it is naturally compatible with heterogeneous LLMs of different structures, as well as most model enhancement methods, including fine-tuning Hu et al. (2022); Zhang et al. (2023b), Mixture-of-Experts Jiang et al. (2024); Zhong et al. (2024a), model merging Goddard et al. (2024); Yang et al. (2024), etc. This allows for the creation of a sufficiently large LLM candidate pool. In this paper, through extensive experiments, we find a model-level scaling up phenomenon in LLMs, i.e., using some capable router, the rapid performance improvement as the number of candidates in the LLM pool increases, can even easily surpass the performance of the best single model in the pool and most existing strong LLMs. Moreover, the majority of LLMs in the candidate pool are relatively small and open-source models whose individual capabilities are far below those of commercialized LLMs, like GPT-4 Achiam et al. (2023), which demonstrates Routing LLM is a highly promising paradigm (See Section 3 for details).
@@
Figure 1: The Overview of Routing LLMs. For each given input, the router distributes it to the appropriate LLM to achieve specific objectives, such as high accuracy, low computational cost, reduced hallucinations, etc. Routing LLMs is a promising paradigm for LLMs to achieve model-level scaling up.
@@
Figure 2: The Model-level Scaling Up Phenomenon in Routing LLMs. 3 As shown in Section 3, the Prob. $p$ indicates the performance of the router, with values closer to 1 representing greater similarity to the oracle router’s capability. If $p\to 0$, then $r_{o}(p)$ degenerates into a random sampler. When the router $r_{o}(p)$ reaches a certain level of capability, it induces a scaling up phenomenon in the Routing LLMs paradigm. Specifically, as the number of LLM candidates increases, performance rapidly improves. "Ref. LLM" denotes a representative LLM with strong performance on given benchmark, such as GPT-4. Further details are provided in Section 5.1. For more examples of model-level scaling up phenomenon, please refer to the Appendix G.
@@
However, current Routing LLMs methods are still in the early stages of rapid development and lack comprehensive and open-source benchmarks specifically designed for routers, which hinders their progress. For example, some existing benchmarks Lu et al. (2024a); Hu et al. (2024a) still have the issues of insufficient LLM candidates, limited evaluation diversity, closed-source or inadequate performance records, etc. To address these challenges, in this paper, we collect and clean 12 popular LLM evaluations covering fields such as knowledge-based Q&A, commonsense inference, semantic understanding, etc. From these evaluations, we use performance records from over 8,500 LLMs, amounting to more than 200,000,000 entries, to construct a comprehensive Router benchmark named RouterEval in Section 4. Utilizing RouterEval, in Section 5, we conduct a comprehensive exploration of several existing router construction methods across various settings and find that these methods still have considerable room for improvement. We show the related works in Appendix B, and summarize the contributions of this paper as follows:
@@ 3 Model-level Scaling Up Phenomenon
[backgroundcolor=gray!8] (2) Weak Candidates Can Also be Promising. In fact, the vast majority of LLMs discussed in this paper are open-source and can be deployed locally (see Appendix F for details). The individual performance of these models is not particularly remarkable and generally falls significantly short of mainstream LLMs. However, as shown in Fig. 2, even relatively weak candidates can achieve complementary performance among multiple heterogeneous models under the Routing LLMs paradigm. They can obtain fine-grained division of labor for each input and outperform mainstream LLMs. This shows that relatively weak candidates can also be promising for obtaining high performance within this paradigm. See more details in Section 4.2.
@@ 4 The Construction of RouterEval
In this section, we elaborate on the construction of our comprehensive benchmark for Routing LLMs, named RouterEval. We collected over 200,000,000 performance records for 12 popular LLM evaluations spanning areas such as knowledge-based Q&A, commonsense reasoning, semantic understanding, and instruction following, based on more than 8,500 LLMs. The benchmarks Zellers et al. (2019); Clark et al. (2018); Wang et al. (2024); Hendrycks et al. (2020); Lin et al. (2021); Sakaguchi et al. (2021); Cobbe et al. (2021); Zhou et al. (2023); Suzgun et al. (2022); Rein et al. (2024); Sprague et al. (2023) involved include ARC, HellaSwag, MMLU, TruthfulQA, Winogrande, GSM8k, IFEval, BBH, GPQA, MUSR, MATH Lvl 5, and MMLU-PRO. See Appendix F, A and C for more details.
@@ 4.2 The Construction of LLM Candidates
When the $m$ LLMs in $G$ are all selected from $\{\ell_{i}^{\prime}\}_{i=1}^{\lfloor 0.2N\rfloor}$ and $\{\ell_{i}^{\prime}\}_{i=\lfloor 0.8N\rfloor}^{N}$ respectively, $\hat{G}$ forms the "all-strong" group and the "all-weak" group. And $\hat{G}$ forms the "strong-to-weak" group, when the $j$-th LLM in $G$ is selected from $\{\ell_{i}^{\prime}\}_{i=(j-1)m}^{\min(jm,N)}$. Through this process, we can explore the router from multiple perspectives and investigate the potential of Routing LLMs. For instance, in MMLU and $m=10$, while the individual performance of each LLM in the "all-weak" group does not exceed 0.3, its oracle performance can reach 0.95. See Appendix C for more details. Additionally, through validation, we ensure that in all candidate groups, at least $10^{30}$ routers can achieve oracle performance, showing the learnability of the routers.
@@ 4.3 Extra Training Data
Table 1: The Results on RouterEval (part1).See Section 5.1 for detials of various metrics. Red area and blue area highlights indicate the "Strong router" and "Existing router" mentioned in Section 5.2, respectively. The best results in existing router methods are highlighted with underlines and on bold. The values in the table are rounded to two decimal places. The results on hard level settings are shown in Appendix E.
@@ 5.2 Baselines
In this paper, we consider two kind of baselines for RouterEval. Specifically, these include: (1) Strong router. To directly evaluate the performance of routers, we introduce the oracle router $r_{o}$ and $r_{o}(0.5)$ mentioned in Section 3 . (2) Existing router. We also compile some recently router methods, such as LinearR, MLPR, C-RoBERTa, MLC, and PRknn Hu et al. (2024a); Srivatsa et al. (2024); Zhao et al. (2024a). See Appendix D or our code for detailed implementation.
@@ 5.3 Result
In this section, we present the results of the baselines shown in Section 5.2, evaluated using the metrics described in Section 5.1. The results under easy level settings, $m\in\{3,5\}$, are shown in Table 1 and Table 2. See more results under hard level settings $m\in\{10,100,1000\}$ in Appendix E.
@@
Table 2: The Results on RouterEval (part2). See Section 5.1 for detials of various metrics. Red area and blue area highlights indicate the "Strong router" and "Existing router" mentioned in Section 5.2, respectively. The best results in existing router methods are highlighted with underlines and on bold. The values in the table are rounded to two decimal places. The results on hard level settings are shown in Appendix E.
@@ 6 Analysis
[backgroundcolor=gray!8] (1) The differences between Routing LLM and existing paradigms. In fact, several existing paradigms in other fields are related to Routing LLMs. This section aims to analyze their relationships and differences. See related works in Appendix B for more details.
@@ Appendix A The More Details of RouterEval
Appendix B Related Work
B.1 Routing LLMs
@@
In contrast, predictive routing leverages machine learning techniques—such as supervised learning (Shnitzer et al., 2023b), reward models (Hari and Thomson, 2023b; Lu et al., 2023b), or meta-models predicting performance scores (Šakota et al., 2023)—to preemptively identify the optimal LLM, thereby reducing latency and computational overhead. Recent advancements have enhanced routing’s effectiveness through innovative techniques, including neural network-based meta-models (Ding et al., 2024a; Šakota et al., 2024; Chen et al., 2024; Aggarwal et al., 2024), k-nearest neighbors (Hu et al., 2024b; Shnitzer et al., 2023c; Stripelis et al., 2024; Lee et al., 2024), matrix factorization (Ong et al., 2024; Zhuang et al., 2024b; Li, 2025), and graph neural networks (Feng et al., 2024). These methods enable coordination among multiple LLMs or even sub-models within a single framework, such as MatFormer (Devvrit et al., 2024; Cai et al., 2024).
@@ B.5 LLM Model Fusion
B.6 Mixture-of-Experts (MoE)

### v2

@@ Abstract
Routing large language models (LLMs) is a new paradigm that uses a router to recommend the best LLM from a pool of candidates for a given input. In this paper, our comprehensive analysis with more than 8,500 LLMs reveals a novel model-level scaling up phenomenon in Routing LLMs, i.e., a capable router can significantly enhance the performance of this paradigm as the number of candidates increases. This improvement can even surpass the performance of the best single model in the pool and many existing strong LLMs, confirming it a highly promising paradigm. However, the lack of comprehensive and open-source benchmarks for Routing LLMs has hindered the development of routers. In this paper, we introduce RouterEval  [^1] See project page for all data, code and tutorial. , a benchmark tailored for router research, which includes over 200,000,000 performance records for 12 popular LLM evaluations across various areas such as commonsense reasoning, semantic understanding, etc., based on over 8,500 various LLMs. Using RouterEval, extensive evaluations of existing Routing LLM methods reveal that most still have significant room for improvement.
@@
Zhongzhan Huang1, Guoming Ling1, Yupei Lin1, Yandong Chen1, Shanshan Zhong1, Hefeng Wu1, Liang Lin1 1Sun Yat-sen University
@@ 1 Introduction
Routing LLMs Jitkrittum et al. (2025); Lu et al. (2023a); Zhao et al. (2024a); Shnitzer et al. (2023a); Chen et al. (2024a) aims to establish an efficient router capable of selecting an LLM from a pool of candidates with varying abilities to effectively handle a given input, in order to achieve specific goals such as high overall accuracy, low computational cost, or minimal hallucination, as shown in Fig. 1. Since this paradigm only involves input assignment, it is naturally compatible with heterogeneous LLMs of different structures, as well as most model enhancement methods, including fine-tuning Hu et al. (2022); Zhang et al. (2023b), Mixture-of-Experts Jiang et al. (2024); Zhong et al. (2024a), model merging Goddard et al. (2024); Yang et al. (2024), etc. This allows for the creation of a sufficiently large LLM candidate pool. In this paper, through extensive experiments, we find a model-level scaling up phenomenon in LLMs, i.e., using some capable router, the rapid performance improvement as the number of candidates in the LLM pool increases, can even easily surpass the performance of the best single model in the pool and most existing strong LLMs. Moreover, the majority of LLMs in the candidate pool are relatively small and open-source models whose individual capabilities are far below those of commercialized LLMs, like GPT-4 Achiam et al. (2023), which demonstrates Routing LLM is a highly promising paradigm (See Section 3 for details).
@@
Figure 1: The Overview of Routing LLMs. For each given input, the router distributes it to the appropriate LLM to achieve specific objectives, such as high accuracy, low computational cost, reduced hallucinations, etc. We find that Routing LLMs is a promising paradigm for LLMs to achieve model-level scaling up.
@@
Figure 2: The Model-level Scaling Up Phenomenon in Routing LLMs. As shown in Section 3, the Prob. $p$ indicates the performance of the router, with values closer to 1 representing greater similarity to the oracle router’s capability. If $p\to 0$, then $r_{o}(p)$ degenerates into a random sampler. When the router $r_{o}(p)$ reaches a certain level of capability, it induces a scaling up phenomenon in the Routing LLMs paradigm. Specifically, as the number of LLM candidates increases, performance rapidly improves. "Ref. LLM" denotes a representative LLM with strong performance on given benchmark, such as GPT-4. Further details are provided in Section 5.1. For more examples of model-level scaling up phenomenon, please refer to the Appendix G.
@@
However, current Routing LLMs methods are still in the early stages of rapid development and lack comprehensive and open-source benchmarks specifically designed for routers, which hinders their progress. For example, some existing benchmarks Lu et al. (2024a); Hu et al. (2024a) still have the issues of insufficient LLM candidates, limited evaluation diversity, closed-source or inadequate performance records, etc. To address these challenges, in this paper, we collect and clean 12 popular LLM evaluations covering fields such as knowledge-based Q&A, commonsense inference, semantic understanding, etc. From these evaluations, we use performance records from over 8,500 LLMs, amounting to more than 200,000,000 entries, to construct a comprehensive Router benchmark named RouterEval in Section 4. Utilizing RouterEval, in Section 5, we conduct a comprehensive exploration of several existing router construction methods across various settings and find that these methods still have considerable room for improvement. We show the related works in Appendix C, and summarize the contributions of this paper as follows:
@@ 3 Model-level Scaling Up Phenomenon
[backgroundcolor=gray!8] (2) Weak Candidates Can Also be Promising. In fact, the vast majority of LLMs discussed in this paper are open-source and can be deployed locally (see Appendix B for details). The individual performance of these models is not particularly remarkable and generally falls significantly short of mainstream LLMs. However, as shown in Fig. 2, even relatively weak candidates can achieve complementary performance among multiple heterogeneous models under the Routing LLMs paradigm. They can obtain fine-grained division of labor for each input and outperform mainstream LLMs. This shows that relatively weak candidates can also be promising for obtaining high performance within this paradigm. See more details in Section 4.2.
@@ 4 The Construction of RouterEval
In this section, we elaborate on the construction of our comprehensive benchmark for Routing LLMs, named RouterEval. We collected over 200,000,000 performance records for 12 popular LLM evaluations spanning areas such as knowledge-based Q&A, commonsense reasoning, semantic understanding, and instruction following, based on more than 8,500 LLMs. The benchmarks Zellers et al. (2019); Clark et al. (2018); Wang et al. (2024); Hendrycks et al. (2020); Lin et al. (2021); Sakaguchi et al. (2021); Cobbe et al. (2021); Zhou et al. (2023); Suzgun et al. (2022); Rein et al. (2024); Sprague et al. (2023) involved include ARC, HellaSwag, MMLU, TruthfulQA, Winogrande, GSM8k, IFEval, BBH, GPQA, MUSR, MATH Lvl 5, and MMLU-PRO. See Appendix B, A and D for more details.
@@ 4.2 The Construction of LLM Candidates
When the $m$ LLMs in $G$ are all selected from $\{\ell_{i}^{\prime}\}_{i=1}^{\lfloor 0.2N\rfloor}$ and $\{\ell_{i}^{\prime}\}_{i=\lfloor 0.8N\rfloor}^{N}$ respectively, $\hat{G}$ forms the "all-strong" group and the "all-weak" group. And $\hat{G}$ forms the "strong-to-weak" group, when the $j$-th LLM in $G$ is selected from $\{\ell_{i}^{\prime}\}_{i=(j-1)m}^{\min(jm,N)}$. Through this process, we can explore the router from multiple perspectives and investigate the potential of Routing LLMs. For instance, in MMLU and $m=10$, while the individual performance of each LLM in the "all-weak" group does not exceed 0.3, its oracle performance can reach 0.95. See Appendix D for more details. Additionally, through validation, we ensure that in all candidate groups, at least $10^{30}$ routers can achieve oracle performance, showing the learnability of the routers.
@@ 4.3 Extra Training Data
Table 1: The Results on RouterEval (part1).See Section 5.1 for detials of various metrics. Red area and blue area highlights indicate the "Strong router" and "Existing router" mentioned in Section 5.2, respectively. The best results in existing router methods are highlighted with underlines and on bold. The values in the table are rounded to two decimal places. The results on hard level settings are shown in Appendix F.
@@ 5.2 Baselines
In this paper, we consider two kind of baselines for RouterEval. Specifically, these include: (1) Strong router. To directly evaluate the performance of routers, we introduce the oracle router $r_{o}$ and $r_{o}(0.5)$ mentioned in Section 3 . (2) Existing router. We also compile some recently router methods, such as LinearR, MLPR, C-RoBERTa, MLC, and PRknn Hu et al. (2024a); Srivatsa et al. (2024); Zhao et al. (2024a). See Appendix E or our code for detailed implementation.
@@ 5.3 Result
In this section, we present the results of the baselines shown in Section 5.2, evaluated using the metrics described in Section 5.1. The results under easy level settings, $m\in\{3,5\}$, are shown in Table 1 and Table 2. See more results under hard level settings $m\in\{10,100,1000\}$ in Appendix F.
@@
Table 2: The Results on RouterEval (part2). See Section 5.1 for detials of various metrics. Red area and blue area highlights indicate the "Strong router" and "Existing router" mentioned in Section 5.2, respectively. The best results in existing router methods are highlighted with underlines and on bold. The values in the table are rounded to two decimal places. The results on hard level settings are shown in Appendix F.
@@ 6 Analysis
[backgroundcolor=gray!8] (1) The differences between Routing LLM and existing paradigms. In fact, several existing paradigms in other fields are related to Routing LLMs. This section aims to analyze their relationships and differences. See related works in Appendix C for more details.
@@ Appendix A The More Details of RouterEval
Appendix B The Details of LLMs Considered
Due to the involvement of different LLMs in constructing performance records for various benchmarks, the LLMs associated with each benchmark vary. Given that each benchmark involves thousands of LLMs and that table space is limited, we are unable to directly display the complete list of all LLMs. Please refer to our code for detailed information. In this section, we provide statistical results of these LLMs for reference.
On the one hand, we specifically show the number of LLMs involved in each benchmark in Table 4. It should be noted that some performance records contain duplicates or errors, and the performance of some open-source models collected is very low (below 0.1), which has almost no reference value. We have carefully filtered these data. It can be seen that the number of LLMs involved in different benchmarks ranges from 1800 to 5000, with a total of more than 8,500 LLMs involved.
Furthermore, we have conducted a statistical analysis of the parameters of all LLMs involved in this study, as shown in Fig. 4. It can be observed that the majority of these LLMs are 7B models, most of which are relatively weak open-source models. Specifically, in Fig. 5, we present the performance distribution of these models across 12 different benchmarks. Nevertheless, the experiments in the main text have demonstrated that even using relatively weak models as candidates, the Routing LLMs paradigm can still fully leverage the strengths of each model, achieve complementary advantages, and ultimately achieve outstanding overall performance.
Figure 4: The Distribution of Model Parameters. We conduct a statistical analysis of the parameter counts for all LLMs considered in this paper and find that models with 7B parameters are predominant.
Figure 5: The Distribution of Model Performance Under 12 Benchmarks. We present the performance distribution of the LLMs involved in each of the 12 benchmarks. As shown in Fig. 4, the majority of LLMs are 7B models, which tend to exhibit relatively weaker performance across the benchmarks from a statistical standpoint.
Appendix C Related Work
C.1 Routing LLMs
@@
In contrast, predictive routing leverages machine learning techniques—such as supervised learning (Shnitzer et al., 2023b), reward models (Hari and Thomson, 2023b; Lu et al., 2023b), or meta-models predicting performance scores (Šakota et al., 2023)—to preemptively identify the optimal LLM, thereby reducing latency and computational overhead. Recent advancements have enhanced routing’s effectiveness through innovative techniques, including neural network-based meta-models (Ding et al., 2024a; Šakota et al., 2024; Chen et al., 2024b; Aggarwal et al., 2024), k-nearest neighbors (Hu et al., 2024b; Shnitzer et al., 2023c; Stripelis et al., 2024; Lee et al., 2024), matrix factorization (Ong et al., 2024; Zhuang et al., 2024b; Li, 2025), and graph neural networks (Feng et al., 2024). These methods enable coordination among multiple LLMs or even sub-models within a single framework, such as MatFormer (Devvrit et al., 2024; Cai et al., 2024).
@@ C.5 LLM Model Fusion
$m$ Group Oracle $r_{o}$ $r_{o}(0.5)$ Details
3 all-strong 0.972 0.924 [0.869, 0.878, 0.879]
3 all_weak 0.765 0.583 [0.406, 0.392, 0.408]
3 strong-to-weak 0.938 0.755 [0.241, 0.599, 0.878]
5 all-strong 0.983 0.930 [0.863, 0.869, 0.878, 0.879, 0.892]
5 all_weak 0.843 0.566 [0.279, 0.300, 0.313, 0.282, 0.276]
5 strong-to-weak 0.943 0.761 [0.330, 0.462, 0.615, 0.622, 0.869]
10 all-strong 0.984 0.881 [0.766, 0.776, 0.781, 0.772, 0.771, 0.783, 0.787, 0.780, 0.778, 0.782]
10 all-weak 0.955 0.620 [0.256, 0.299, 0.276, 0.296, 0.284, 0.287, 0.274, 0.280, 0.301, 0.278]
10 strong-to-weak 0.981 0.765 [0.245, 0.259, 0.487, 0.534, 0.577, 0.614, 0.606, 0.628, 0.658, 0.869]
100 all-strong 0.996 0.890 See Fig. 6
100 all-weak 1.000 0.630 See Fig. 6
100 strong-to-weak 1.000 0.769 See Fig. 6
1000 all-strong 1.000 0.839 See Fig. 6
1000 all-weak 1.000 0.642 See Fig. 6
1000 strong-to-weak 1.000 0.769 See Fig. 6
Table 5: The Details of Candidate Groups on MMLU. $m$ denotes the number of LLMs per candidate group. "Details" shows the performance of each LLM in the corresponding group. When $m=100$ or $m=1000$, the number of candidates is too large to present individually. Therefore, we only display their performance distributions in Fig. 6. For specific performance metrics and the details of other evaluations, please refer directly to our code.
Figure 6: The Distribution of Model Performance in Candidate Group on MMLU. Due to the space limitations of Table 5, it is not feasible to present the detailed performance of all candidates in the candidate groups where $m=100$ or $m=1000$. Therefore, we illustrate their performance distributions in this figure. For specific performance metrics and the details of other evaluations, please refer directly to our code.
C.6 Mixture-of-Experts (MoE)